# Interpretation of the Boundary Current Synchronization as a Maxwell’s Demon

### v1

@@ Abstract
This study has applied information thermodynamics to a bivariate linear stochastic differential equation (SDE) that describes a synchronization phenomenon of sea surface temperatures (SSTs) between the Gulf Stream and the Kuroshio Current, which is referred to as the boundary current synchronization (BCS). Information thermodynamics divides the entire system fluctuating with stochastic noise into subsystems and describes the interactions between these subsystems from the perspective of information transfer. The SDE coefficients have been estimated through regression analysis using observational and numerical simulation data. In the absence of stochastic noise, the solution of the estimated SDE shows that the SSTs relax toward zero without oscillating. The estimated SDE can be interpreted as a Maxwell’s demon system, with the Gulf Stream playing the role of the “Particle” and the Kuroshio Current playing the role of the “Demon.” This interpretation gives the asymmetric roles of both ocean currents. The Gulf Stream forces the SST of the Kuroshio Current to be in phase. By contrast, the Kuroshio Current maintains the phase by interfering with the relaxation of the Gulf Stream SST. In the framework of Maxwell’s demon, the Gulf Stream is interpreted as being measured by the Kuroshio Current, whereas the Kuroshio Current is interpreted as performing feedback control on the Gulf Stream. When the Gulf Stream and the Kuroshio Current are coupled in an appropriate parameter regime, synchronization is realized with atmospheric and oceanic noise as the driving source.
@@ 1 Introduction
Kohyama et al. (2021) discovered the boundary current synchronization (BCS) by using observational data and numerical experiments with coupled atmosphere-ocean models. The sea surface temperatures (SSTs) of the Gulf Stream and the Kuroshio Current show positive correlations on interannual to decadal time scales. An increase (decrease) in the SSTs is accompanied by a northward (southward) shift of the WBCs as well as a northward (southward) shift of the atmospheric baroclinic jet. The BCS is considered to arise from the interaction between the Gulf Stream and the Kuroshio Current via the atmosphere, namely the inter-basin coupling, rather than a passive response to atmospheric forcings. In fact, the BCS was not reproduced in low-resolution numerical experiments that could not resolve oceanic mesoscale eddies. These results suggest that the BCS occurs through the transfer of information on WBC variability through the atmosphere. However, the roles of the Gulf Stream and the Kuroshio Current in the BCS have not been well investigated, and further research is required.
@@ 2 Review of Maxwell’s demon
The resolution of Maxwell’s demon is a landmark in information thermodynamics (Parrondo et al., 2015; Peliti and Pigolotti, 2021; Shiraishi, 2023). By considering the exchange of information between subsystems, this paradox is theoretically resolved (Sagawa and Ueda, 2008, 2009, 2010, 2012). Here, we review the resolution of Maxwell’s demon using a simple model called the Szilard engine. The following discussion is based on a textbook by Sagawa (2022).
2.1 Szilard engine
@@
Let us consider a single particle in a box and construct the following isothermal cycle. The box is surrounded by an isothermal environment, where heat is transferred through the walls of the box. Figure 1 is a schematic of the experimental setup, where the memory (i.e., the demon) in the upper half is not considered for the moment. First, a partition is inserted in the middle of the box, separating the box into halves (Fig. 1a). The particle moves at random, which means that the probability of the particle being on the left (or right) side equals $1/2$. Second, the position of the particle is measured (Fig. 1b). We assume that this measurement does not require work. Third, based on the measurement outcome, only when the particle is on the right (i.e., the lower case in Fig. 1), the right half of the box is moved quasi-statically to the left (Fig. 1c). During this process, the forces are balanced; thus, no work is necessary. Such a measurement-based operation is called feedback control or simply, control. After this operation, the particle is on the left with a probability of $1$. Finally, the wall on the right side is moved to the right, the box is expanded isothermally (Fig. 1d), and the particle system returns to the initial state. The energy source for the expansion is the surrounding thermal fluctuations, or equivalently, heat transferred through the box. Therefore, this cycle extracts work from the isothermal environment. The Szilard engine is apparently a perpetual motion machine of the second kind and contradicts the second law of thermodynamics that is expressed as the Kelvin statement.
@@
Figure 1: Schematic of the Szilard engine, a simple model illustrating Maxwell’s demon (Sagawa, 2022). A box contains a single particle and is surrounded by an isothermal environment, where heat is transferred through the walls of the box. (a) A partition is inserted in the middle of the box. (b) The demon measures the position of the particle and records the result in its memory. (c) Based on the measurement outcome, the demon performs feedback control. Specifically, only when the particle is on the right (i.e., the lower case in the figure), the right half of the box is moved quasi-statically to the left. (d) The demon expands the box isothermally and returns the particle system to the initial state. This process extracts work from the isothermal environment (i.e., a positive work). (e) The demon’s memory is reset to its initial state, requiring external work (i.e., a negative work).
@@
To resolve the paradox, it is necessary to consider the exchange of information quantities and the existence of memory (Sagawa and Ueda, 2013a; Sagawa, 2022). The information quantity is precisely called mutual information (e.g., Cover and Thomas, 2005), which is an indicator of the nonlinear correlation between subsystems. The memory, also called the demon, is a subsystem that performs measurement and feedback control. The other subsystem is the particle within the box and is referred to as the particle. We denote the states of the demon and the particle by $X$ and $Y$ and their realizations by lowercase letters $x$ and $y$, respectively. The variables $X$ and $Y$ are discrete, reflecting the particle position being either on the left or right. We first derive the second law of information thermodynamics and then consider the Szilard engine again. More detailed discussions can be found in textbooks (Peliti and Pigolotti, 2021; Shiraishi, 2023).
@@ 2.2 The second law of information thermodynamics
We define Shannon entropy and mutual information to derive the second law of information thermodynamics.
@@
Shannon entropies $S(X)$ and $S(Y)$ are indicators of uncertainty about the subsystems $X$ and $Y$, respectively (e.g., Cover and Thomas, 2005). Mutual information $M(X,Y)$ is a non-negative indicator of the correlation between $X$ and $Y$ (e.g., Cover and Thomas, 2005), where $M(X,Y)$ is zero only when $X$ and $Y$ are statistically independent. Following the convention, we distinguish probability distributions by their arguments. For instance, $p(x)$ and $p(y)$ are probability distributions for the random variables $X$ and $Y$, respectively, and are different from each other.
@@
The second law of information thermodynamics can be derived using the non-additivity of Shannon entropy (Sagawa and Ueda, 2013b). We consider the evolution from time $t$ to $t+\Delta t$ and assume the second law of thermodynamics for the entire system (Seifert, 2005, 2012):
@@
where changes are denoted by $\Delta$, and time dependence is denoted by a subscript $t$. For example, $\Delta S(X_{t},Y_{t}):=S(X_{t+\Delta t},Y_{t+\Delta t})-S(X_{t},Y_{t})$. The symbol $\Delta Q_{t}$ represents the amount of heat dissipated from the joint system $(X,Y)$ to the surrounding environment. This environment is sufficiently large compared to the system $(X,Y)$ and always in a thermal equilibrium state at temperature $T$. In this case, the entropy change of the environment is expressed as $\Delta Q_{t}/T$. Equation (6), namely the second law of thermodynamics, expresses that the total entropy change, including the environment, is non-negative.
The second law of information thermodynamics for the subsystem $X$ can be derived from Eq. (6) by assuming that only $X$ changes and $Y$ does not change in time. Using the entropy decomposition [Eq. (5)], the following is derived:
@@
where $\Delta S(Y)=0$ because $Y_{t}$ does not change in time. This assumption also indicates that the entropy change of the environment, $\Delta Q_{t}/T$, is due to the change in $X_{t}$. We finally obtain the second law of information thermodynamics:
@@
The total entropy change by $X_{t}$ is denoted as $\Delta\sigma(X_{t})$ and is lower bounded by the change in mutual information $\Delta M\left(X_{t},Y_{t}\right)$. Since $Y_{t}$ does not change, $\Delta M\left(X_{t},Y_{t}\right)$ is attributed only to the change in $X_{t}$. The second law of information thermodynamics indicates that $\Delta\sigma(X_{t})$ can be negative when $\Delta M\left(X_{t},Y_{t}\right)<0$. Thus, the entropy of the subsystem $X$ can be decreased by utilizing the change in mutual information without heat transfer. This conclusion is consistent with the second law of thermodynamics because this discussion started from the second law of thermodynamics for the entire system [Eq. (6)]. Note that $\Delta\sigma(Y_{t})\geq M(X_{t},Y_{t+\Delta t})-M(X_{t},Y_{t})$ is derived when $X$ does not change and only $Y$ changes in time.
2.3 Resolution of Maxwell’s demon
@@
The second law of information thermodynamics is used to resolve Maxwell’s demon paradox. Time $t$ is not explicitly stated if there is no confusion. In the initial state, the states of the demon $X$ and the particle $Y$ are uncorrelated, that is, $M(X,Y)$ is zero. First, the demon inserts a partition (Fig. 1a). This causes the particle $Y$ to be either on the right or the left. The mutual information $M(X,Y)$ remains zero because the demon’s internal state (i.e., the memory $X$) has not changed.
@@
Next, the demon measures the position of the particle and changes its memory $X$ to record the measurement outcome, which makes $X$ either right or left (Fig. 1b). The states of the demon and the particle become correlated, and the mutual information increases, $\Delta M(X,Y)>0$. Since the particle $Y$ does not change during this measurement process, this increase in correlation is due to the change in the demon $X$. Therefore, the demon is considered to gain mutual information. According to the second law of information thermodynamics, this information gain is accompanied by a positive entropy change, $\Delta\sigma(X)\geq\Delta M(X,Y)$ ($>0$) [Eq. (8)].
@@
Based on the memory value $X$, the demon quasi-statically moves the box, only when the particle is on the right (Fig. 1c). The particle is then on the left with a probability of $1$. The demon expands the box isothermally and returns the particle $Y$ to the initial state (Fig. 1d). Regardless of the memory value, the particle $Y$ returns to the initial state after these feedback control processes, and the correlation $M(X,Y)$ becomes zero again. Since the demon’s memory $X$ does not change during these processes, this decrease in correlation is due to the change in the particle $Y$. Therefore, the particle is considered to lose mutual information. According to the second law of information thermodynamics, this information loss can be accompanied by a negative entropy change, that is, $\Delta\sigma(Y)\geq\Delta M(X,Y)$ ($<0$). The negative entropy change $\Delta\sigma(Y)$ corresponds to the positive work extracted by the isothermal expansion.
@@
Finally, work is performed from the outside to return the demon’s memory $X$ to the initial state (Fig. 1e). For simplicity, we have assumed that no work is required for the measurement (Section 22.1). There is a trade-off for the memory between initialization and measurement, and it is more appropriate to consider both together (Sagawa and Ueda, 2009, 2013b; Sagawa, 2019). The positive work extracted by the isothermal expansion and the negative work required for the initialization (and measurement) cancel each other out. Thus, no net positive work can be extracted using the Szilard engine. Therefore, the Szilard engine is not a perpetual motion machine of the second kind.
@@
Information thermodynamics clarifies that the second law of thermodynamics [Eq. (6)] is not violated in each process of measurement and control because Eq. (8) was derived from Eq. (6). In particular, the positive work extracted during the control process is due to the change in mutual information. This means that the total entropy of a subsystem can decrease by consuming mutual information. This kind of entropy decrease is newly revealed by information thermodynamics because usual thermodynamics shows that the total entropy of a subsystem can be decreased only through heat exchange. In this way, information thermodynamics allows for the analysis of subsystems from the perspective of information.
2.4 Maxwell’s demon in autonomous systems
@@
Maxwell’s demon is also realized in autonomous systems (Horowitz and Esposito, 2014; Loos and Klapp, 2020). The time evolution of an autonomous system can be expressed by stochastic differential equations without external interference, such as human manipulation. In such an autonomous system, $X$ and $Y$ change continuously in time, and it is difficult to stop the change of one and examine the other’s time evolution. In other words, for autonomous systems, the distinction between the “Demon” and the “Particle” is not necessarily clear.
@@
The determination of which subsystem is the “Demon” or the “Particle” is based on the direction of information flow (Allahverdyan et al., 2009; Horowitz and Esposito, 2014; Loos and Klapp, 2020). The information flow is the rate of change of mutual information due to the change in either $X$ or $Y$.
@@
The information flow $\dot{I}_{X\leftarrow Y}$ from $Y$ to $X$ is the change in mutual information $M(X_{t},Y_{t})$ due to the change in $X$, whereas $\dot{I}_{Y\leftarrow X}$ is the change in $M(X,Y)$ due to the change in $Y$. The sum of both information flows is equal to the total rate of change of mutual information [Eq. (13)]. Information flow is regarded as a quantity that represents the flow of information between subsystems. For instance, when $\dot{I}_{X\leftarrow Y}$ is positive (negative), the subsystem $X$ is considered to gain (lose) information from the subsystem $Y$.
For autonomous systems that evolve continuously in time, the second law of information thermodynamics becomes as follows (Horowitz and Esposito, 2014; Loos and Klapp, 2020):
@@
The rate of change of total entropy is lower bounded by the information flow. Hereafter, the rate of change of total entropy is referred to as the production rate of total entropy or simply the entropy production rate. Section 44.1 gives the expressions for the information flows and the production rates of total entropy. The equations used there are specific forms for a bivariate linear dynamical system introduced in Section 3. A general theory of information thermodynamics, including nonlinearity, is reviewed in Appendix A.
In general, the production rate of total entropy is an indicator of irreversibility. For a subsystem, when the production rate of total entropy equals the information flow, that is, when the equality holds in the second law of information thermodynamics, the change in the subsystem becomes reversible (Horowitz and Esposito, 2014; Loos and Klapp, 2020; Seifert, 2012).
However, this study does not use the connection between the second law and irreversibility. The second law of information thermodynamics generally holds for a system described by stochastic differential equations (Appendix A) and thus can be regarded as a constraint on the time evolution of the system. In other words, information thermodynamics can be employed as a mathematical tool to analyze the time evolution of the system. Following this approach, we analyze the dynamics of an idealized system for the BCS by relating the entropy production and information flow to the system’s dynamics (Section 4).
We finally summarize the distinction between a “Demon” and a “Particle.” As in Section 22.3, measurement is a change that increases correlation due to an evolution of the “Demon”, which is accompanied by a positive entropy production. On the other hand, control is a change that decreases correlation due to an evolution of the “Particle”, which is accompanied by a negative entropy production. Therefore, in an autonomous system, the “Demon” is the subsystem showing a positive information flow and a positive production rate of total entropy, whereas the “Particle” is the subsystem showing a negative information flow and a negative production rate of total entropy (Horowitz and Esposito, 2014; Loos and Klapp, 2020). In this study, a system consisting of a “Demon” and a “Particle” is referred to as the Maxwell’s demon system. Maxwell’s demon systems are no longer thought experiments but have been experimentally realized in both non-autonomous and autonomous settings (Toyabe et al., 2010; Koski et al., 2015). We propose that such systems can also be realized in the atmosphere and ocean, that is, the coupled system of the Gulf Stream and the Kuroshio Current can be interpreted as a Maxwell’s demon system.

### v2

# A Mechanism of Stochastic Synchronization in the Climate System: an Interpretation of the Boundary Current Synchronization as a Maxwell’s Demon

@@ Abstract
This study has applied information thermodynamics to a bivariate linear stochastic differential equation (SDE) that describes a synchronization phenomenon of sea surface temperatures (SSTs) between the Gulf Stream and the Kuroshio Current, which is referred to as the boundary current synchronization (BCS). Information thermodynamics divides the entire system fluctuating with stochastic noise into subsystems and describes the interactions between these subsystems from the perspective of information transfer. The SDE coefficients have been estimated through regression analysis using observational and numerical simulation data. In the absence of stochastic noise, the solution of the estimated SDE shows that the SSTs relax toward zero without oscillation. The estimated SDE can be interpreted as a Maxwell’s demon system, with the Gulf Stream playing the role of the “Particle” and the Kuroshio Current playing the role of the “Demon.” The Gulf Stream forces the SST of the Kuroshio Current to be in phase. By contrast, the Kuroshio Current maintains the phase by interfering with the relaxation of the Gulf Stream SST. In the framework of Maxwell’s demon, the Gulf Stream is measured by the Kuroshio Current, whereas the Kuroshio Current performs feedback control on the Gulf Stream. When the Gulf Stream and the Kuroshio Current are coupled in an appropriate parameter regime, synchronization is realized with atmospheric and oceanic fluctuations as the driving source. This new mechanism, “stochastic synchronization,” suggests that such fluctuations can be converted into directional variations in a subsystem of the climate system through synchronization by utilizing information (i.e., information-to-energy conversion).
@@ 1 Introduction
Stochastic modeling is an effective approach for various climate phenomena (e.g., Dijkstra, 2013; Franzke et al., 2015). This approach is based on the time scale separation between slow and fast variations (Hasselmann, 1976). Typically, slow processes include decadal climate variations, while fast processes include weather disturbances. The scale separation suggests that the correlation between slow and fast processes is weak; thus, the fast processes can effectively be represented by stochastic noise. Stochastic modeling describes the integral (or aggregated) responses of slow climate variables driven by stochastic noise. A major approach in stochastic modeling is the use of idealized dynamical systems with few variables (e.g., Dijkstra, 2013). Although these models cannot describe phenomena in detail, they are expected to capture the essence of phenomena (e.g., Benzi, 2010). In this study, we investigate the coupling of two western boundary currents using a stochastic dynamical system and propose a new mechanism for synchronization, which may be applicable to other climate phenomena.
@@
Kohyama et al. (2021) discovered the boundary current synchronization (BCS) by using observational data and numerical experiments with coupled atmosphere-ocean models. The sea surface temperatures (SSTs) of the Gulf Stream and the Kuroshio Current show positive correlations on interannual to decadal time scales. An increase (decrease) in the SSTs is accompanied by a northward (southward) shift of the WBCs as well as a northward (southward) shift of the atmospheric baroclinic jet. The BCS is considered to arise from the interaction between the Gulf Stream and the Kuroshio Current via the atmosphere (i.e., the inter-basin coupling) rather than a passive response to atmospheric forcings. In fact, the BCS was not reproduced in low-resolution numerical experiments that could not resolve oceanic mesoscale eddies. These results suggest that the BCS occurs through the transfer of information on WBC variability through the atmosphere. However, the roles of the Gulf Stream and the Kuroshio Current in the BCS have not been well investigated, and further research is required.
@@ 2 Review of Maxwell’s demon
The resolution of Maxwell’s demon is a landmark in information thermodynamics (Parrondo et al., 2015). Originally, Maxwell argued that if an intelligent being (a demon) had information about the velocities of particles, it could separate fast, hot particles from slow, cold particles without exerting any work on the system, thus appearing to violate the second law of thermodynamics. Szilard then developed a simple model, the Szilard engine, which captures the essential point of Maxwell’s demon: information can be converted to work. Nowadays, this information-to-energy conversion has been realized experimentally (Toyabe et al., 2010; Koski et al., 2014). This type of system is called an information engine or loosely Maxwell’s demon (e.g., Horowitz and Esposito, 2014; Ito and Sagawa, 2015). Here, we refer to the entire system of information conversion as the Maxwell’s demon system and review the Szilard engine as such a system, which consists of a demon (i.e., a controller) and a subsystem operated by the demon. See further details in reviews and textbooks (Sagawa and Ueda, 2013a; Parrondo et al., 2015; Peliti and Pigolotti, 2021; Shiraishi, 2023).
2.1 Setup of the Szilard engine
@@
Let us consider an isothermal cycle with a single particle. Figure 1 is a schematic of this experiment, where the memory (i.e., the demon) in the upper half is not considered for the moment. First, a partition is inserted in the middle of the box, separating the box into halves (Fig. 1a). The particle moves at random, so the probability of the particle being on the left (or right) side equals $1/2$. Second, the position of the particle is measured (Fig. 1b). Third, based on the measurement outcome, when the particle is on the left (resp. right), the partition is moved quasi-statically to the right (resp. left). (Fig. 1c). Such a measurement-based operation is called feedback control or simply, control. The energy source for this expansion is the surrounding thermal fluctuations, or equivalently, heat transferred through the box. Therefore, the Szilard engine appears to contradict the second law of thermodynamics and to be a perpetual motion machine of the second kind that extracts work via isothermal cycles.
@@
Figure 1: Schematic of the Szilard engine (e.g., Parrondo et al., 2015). A box contains a single particle and is surrounded by an isothermal environment, where heat is transferred through the walls of the box. (a) A partition is inserted in the middle of the box. (b) The demon measures the position of the particle and records the result in its memory. (c) Based on the measurement outcome, the demon performs feedback control. Specifically, when the particle is on the left (resp. right), the demon moves the partition to the right (resp. left). The particle system is isothermally expanded and returns to the initial state. This process extracts a positive work from the isothermal environment. (d) The demon’s memory is initialized, requiring external work (i.e., a negative work).
@@
To resolve the paradox, it is necessary to consider the exchange of information quantities and the existence of memory (Sagawa and Ueda, 2013a). The memory, also called the demon, is a subsystem that performs measurement and feedback control. The other subsystem is the particle within the box and is referred to as the particle. We denote the states of the demon and the particle by $X$ and $Y$ and their realizations by lowercase letters $x$ and $y$, respectively. The variables $X$ and $Y$ are discrete, reflecting the particle position being either on the left or right. We first derive the second law of information thermodynamics and then consider the Szilard engine again.
@@ 2.2 The second law of information thermodynamics
We define Shannon entropy and mutual information (e.g., Cover and Thomas, 2005) to derive the second law of information thermodynamics.
@@
Shannon entropies $S(X)$ and $S(Y)$ are indicators of uncertainty about the subsystems $X$ and $Y$, respectively. Mutual information $M(X,Y)$ is a non-negative indicator of the correlation between $X$ and $Y$, where $M(X,Y)$ is zero only when $X$ and $Y$ are statistically independent.
@@
The second law of information thermodynamics is derived using the non-additivity of Shannon entropy (Sagawa and Ueda, 2013b). We consider the evolution from time $t$ to $t+\Delta t$ and assume the second law of thermodynamics for the entire system (Seifert, 2005, 2012):
@@
where changes are denoted by $\Delta$, and time dependence is denoted by a subscript $t$. For example, $\Delta S(X_{t},Y_{t}):=S(X_{t+\Delta t},Y_{t+\Delta t})-S(X_{t},Y_{t})$. The symbol $\Delta Q_{t}$ represents the amount of heat dissipated from the joint system $(X,Y)$ to the surrounding environment, which corresponds to the entropy change of the environment. Equation (6) expresses that the total entropy change, including the environment, is non-negative.
The second law of information thermodynamics for the subsystem $X$ is derived from Eq. (6) by assuming that only $X$ changes in time. Using the entropy decomposition [Eq. (5)], the following is derived:
@@
where $\Delta S(Y)=0$ because $Y_{t}$ does not change in time. This assumption also indicates that the entropy change of the environment $\Delta Q_{t}/T$ is due to the change in $X_{t}$. We finally obtain the following inequality:
@@
The total entropy change by $X_{t}$ is denoted as $\Delta\sigma(X_{t})$ and is lower bounded by the change in mutual information $\Delta M\left(X_{t},Y_{t}\right)$. Since $Y_{t}$ does not change, $\Delta M\left(X_{t},Y_{t}\right)$ is attributed only to the change in $X_{t}$. The second law of information thermodynamics [Eq. (8)] indicates that $\Delta\sigma(X_{t})$ can be negative when $\Delta M\left(X_{t},Y_{t}\right)<0$. Thus, the entropy of the subsystem $X$ can be decreased by utilizing the change in mutual information without heat transfer. This conclusion is consistent with thermodynamics because this discussion started from the second law of thermodynamics for the entire system [Eq. (6)]. Note that $\Delta\sigma(Y_{t})\geq M(X_{t},Y_{t+\Delta t})-M(X_{t},Y_{t})$ is derived when only $Y$ changes in time.
2.3 Explanation of the Szilard engine
@@
The second law of information thermodynamics is used to explain the Szilard engine. Time $t$ is not explicitly stated if there is no confusion. In the initial state, the states of the demon $X$ and the particle $Y$ are uncorrelated, that is, $M(X,Y)$ is zero. First, the demon inserts a partition (Fig. 1a). This causes the particle $Y$ to be either on the right or the left. The mutual information $M(X,Y)$ remains zero because the demon’s internal state (i.e., the memory $X$) has not changed.
@@
Next, the demon measures the position of the particle and changes its memory $X$ to record the measurement outcome, which makes $X$ either right or left (Fig. 1b). The states of the demon and the particle become correlated, and the mutual information increases, $\Delta M(X,Y)>0$. This increase in correlation is due to the change in the demon $X$. According to the second law of information thermodynamics, the increase in correlation is accompanied by a positive entropy change, $\Delta\sigma(X)\geq\Delta M(X,Y)$ ($>0$) [Eq. (8)].
@@
Based on the memory value $X$, the demon quasi-statically moves the partition to extract work (Fig. 1c). Regardless of the memory value, the particle $Y$ returns to the initial state after this feedback control, and the correlation $M(X,Y)$ becomes zero again. This decrease in correlation is due to the change in the particle $Y$. According to the second law of information thermodynamics, the decrease in correlation can be accompanied by a negative entropy change, $\Delta\sigma(Y)\geq\Delta M(X,Y)$ ($<0$). This negative $\Delta\sigma(Y)$ corresponds to the positive work extracted by the isothermal expansion.
@@
Finally, work is performed from the outside to initialize the demon’s memory $X$ (Fig. 1d). There is a trade-off for the memory between initialization and measurement, and it is more appropriate to consider both together (Sagawa and Ueda, 2009, 2013b; Sagawa, 2019). The positive work extracted by the isothermal expansion and the negative work required for the initialization (and measurement) cancel each other out. No net positive work can be extracted using the Szilard engine. Therefore, the Szilard engine is not a perpetual motion machine of the second kind.
@@
Information thermodynamics clarifies that the positive work extracted during the control process is due to the change in mutual information; that is, the total entropy of a subsystem can decrease by consuming mutual information. In this way, information thermodynamics allows for the analysis of subsystems from the perspective of information.
2.4 Autonomous Maxwell’s demon systems
@@
Maxwell’s demon systems are also realized in an autonomous manner (Horowitz and Esposito, 2014; Loos and Klapp, 2020). The time evolution of an autonomous system can be expressed by stochastic differential equations without external interference, such as human manipulation. In such a system, one cannot isolate individual steps, such as measurement and control, since all steps can occur simultaneously and continuously over time. This is a major difference between the Szilard engine and autonomous processes. Thus, the distinction between the “Demon” and the “Particle” appears unclear.
@@
The determination of which subsystem is the “Demon” or the “Particle” is based on the direction of information flow (Allahverdyan et al., 2009; Horowitz and Esposito, 2014; Loos and Klapp, 2020).
@@
The information flow $\dot{I}_{X\leftarrow Y}$ from $Y$ to $X$ is the change in mutual information $M(X_{t},Y_{t})$ due to the change in $X$, whereas $\dot{I}_{Y\leftarrow X}$ is due to the change in $Y$. The sum of both information flows is equal to the total rate of change of mutual information [Eq. (13)]. Information flow is regarded as a quantity that represents the flow of information between subsystems.
For autonomous systems, the second law of information thermodynamics becomes as follows (Horowitz and Esposito, 2014; Loos and Klapp, 2020):
@@
The rate of change of total entropy is lower bounded by the information flow. Hereafter, the rate of change of total entropy is referred to as the production rate of total entropy or simply the entropy production rate. Equations (14) and (15) are derived for general autonomous systems (Appendix A).
We finally summarize the distinction between a “Demon” and a “Particle.” As in Section 22.3, measurement is a change that increases correlation due to an evolution of the “Demon”, which is accompanied by a positive entropy production. On the other hand, control is a change that decreases correlation due to an evolution of the “Particle”, which is accompanied by a negative entropy production. Therefore, in an autonomous system, the “Demon” (resp. “Particle”) is the subsystem showing a positive (resp. negative) information flow and entropy production rate (Horowitz and Esposito, 2014; Loos and Klapp, 2020). In our study, a system consisting of a “Demon” and a “Particle” has been referred to as the Maxwell’s demon system. Maxwell’s demon systems are no longer thought experiments but have been experimentally realized in both non-autonomous and autonomous settings (Toyabe et al., 2010; Koski et al., 2015). We propose that such systems can also be realized in the atmosphere and ocean, that is, the coupled system of the Gulf Stream and the Kuroshio Current can be interpreted as a Maxwell’s demon system.